If m<1+m<2+m<3= 180 degree, what is the measure of <4?
Answer by chessace(471)

If m is a multiply factor, and the object is a quadrilateral, then
<1+<2+<3+<4 = 360
<4 = 360 -(<1+<2+<3)
But m(<1+<2+<3) = 180
So (<1+<2+<3) = 180/m
Answer: <4 = 360 - 180/m
Note: 1/2 < m < 1.